4. (2pt) Exam Grades Which of these exam grades has a better relative position? a. A grade of 82 on a test with X- 85 and s- 6 b. A grade of 56 on a test with X 60 and s 5 (4pt) ROUND THE ANSWER TO THREE DECIMAL PLACES. Paper Use Each American uses an average of 650 pounds (295 kg) of paper in a year. Suppose that the distribution is approximately normal with a population standard deviation of 153.5 pounds. Assume the variable is normally distributed. Find the probability that a randomly selected American uses a. More than 800 pounds of paper in a year 5. YOU MUST SHOW AT LEAST 2 CURVES
Answer #1

(4)

(a) z = (x - μ)/σ = (82 - 85)/6 = -0.5 for the first exam

z = (x - μ)/σ = (56 - 60)/5 = -0.8 for the second exam

Since -0.5 > -0.8, the grad on exam 1 has a better relative position

(5)

μ = 650, σ = 153.5, x = 800

z = (x - μ)/σ = (800 - 650)/153.5 = 0.9772

P(x > 800) = P(z > 0.9772) = 0.1642